A cable hangs in a parabolic arc between two poles 100 feet apart. The poles are 30 feet high and the lowest point on a suspended cable is 5 feet above the ground.

a) find the equation of the arc if the vertex is the lowest point of the cable

b) the opening of a cave 20 feet wide is in the shape of a semi ellipse. If the height of the opening is 15 feet at the center, how high is the opening 5 feet from the center.



1
Expert's answer
2021-11-10T18:24:56-0500

a) Let point "(0, 5)" be the vertex of the parabola. Then the equation of the parabola will be


"y=ax^2+5, a>0"

Given "y(-50)=y(50)=30." Substitute


"a(50)^2+5=30"

"a=0.01"

The equation of the arc is

"y=0.01x^2+5"

b)

The equation of the ellipse is


"\dfrac{(x-h)^2}{a^2}+\dfrac{(y-k)^2}{b^2}=1"

Substitute


"\dfrac{(x-0)^2}{(20/2)^2}+\dfrac{(y-0)^2}{15^2}=1"

The equation of the given semi ellipse is

"\dfrac{x^2}{100}+\dfrac{y^2}{225}=1, y\geq0"

If "x_1=5"


"\dfrac{(5)^2}{100}+\dfrac{y^2}{225}=1, y\geq0"

"y=\sqrt{225(1-1/4)}"

"y=7.5\sqrt{3}"

The opening 5 feet from the center is "7.5\sqrt{3}(\approx13)" feet high.